If a woman who is a carrier for Fragile X syndrome (an X-linked disorder) has children with a man who does not have the disorder, what chance do their daughters have of being carriers of Fragile X? none 50% 100% 25%
Answer #1

Answer-

Probability of having daughter is 1/2 or 50%

Probability of having daughter who is carrier of Fragile X syndrome is 1/2 \times 1/2 = 1/4 or 25%

Calculation is given below:

Cenotype of caocier croman for fragile x syndrome Genotype of normal man = XOY xDxd x toy 글살금 carrier daughters genotype = x